Using the given equation find the missing coordinates of the points and then find the slope of the line for each equation
y=2x+3; A(...; 5), B(10; ...)
A (... 5), B (10...)
The slope is..

Answer by ankor@dixie-net.com(22740)

Using the given equation find the missing coordinates of the points and then find the slope of the line for each equation
y = 2x + 3; A(...; 5), B(10; ...)
find x when y = 5
2x + 3 = 5
2x = 5 - 3
2x = 2
x = 1
Find y when x = 10
y = 2(10) + 3
y = 23
:
A (1, 5), B (10, 23)
The slope is 2